# Can NeRFs See without Cameras?

### v2

@@ 5 Follow ups and Conclusion
“Can NeRFs See without Cameras?”
@@ Appendix D Relaxing Tx Assumptions
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ E.4 Heatmap Segmentation Implementation Details
Code: We plan to release our code, data, and baselines upon acceptance of the paper to ensure reproducibility. In the meantime, Section E provides sufficient details to allow readers to reproduce our results, especially since our network components are simple MLPs.
@@ Appendix F Evaluation on Additional Floorplans
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption

### v3

@@ Acknowledgments
Can NeRFs “See” without Cameras?
@@ Appendix D Relaxing Tx Assumptions
6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ E.4 Heatmap Segmentation Implementation Details
Code: We plan to release our code, data, and baselines soon. In the meantime, Section E provides sufficient details to allow readers to reproduce our results, especially since our network components are simple MLPs.
@@ Appendix F Evaluation on Additional Floorplans
6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6cm. EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6cm. EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption